OpenStudy (anonymous):

it took Anthony 5 h 45 min to complete a 25-mi walkathon. What was his average speed in miles per hour? Round your answer to the nearest hundredth. a.about 4.35 mi/h b.about 2.27 mi/h c.about 3.34 mi/h d.about 1.55 mi/h

OpenStudy (anonymous):

sumbul comment so i can give you a medal to

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i still need the answer

OpenStudy (anonymous):

it is a

OpenStudy (anonymous):

thanks\

OpenStudy (anonymous):

5 hour 45 minutes =5+(45/60) hour=5+.75=5.75 hour and then divide 25 by 5.75

OpenStudy (lilg132):

a. 4.347 mi/h 5 hours 45 = 345 minutes 345 / 25 = 13.8 minutes per mile 60 / 13.8 = 4.347 mi/h

OpenStudy (anonymous):

both methods are equivalent

OpenStudy (anonymous):

either convert from hour to min or min to hour

OpenStudy (anonymous):

ok thanks

OpenStudy (anonymous):

From Mathematica:\[\text{Round}\left[\frac{25}{\left(5+\frac{45}{60}\right)},0.01\right]=4.35\text{mph} \]